OpenStudy (anonymous):

In a college, some courses contribute more towards an overall GPA than other courses. For example, a science class is worth 4 points; mathematics is worth 3 points; history is worth 2 points; and English is worth 3 points. The values of the grade letters are as follows, A= 4, B=3, C=2, D=1, F=0. What is the GPA of a student who made a “C” in Trigonometry, a “B” in American History, an “A” in Botany, and a “B” in Microbiology? A. 2.59 B. 2.86 C. 3.08 D. 3.33 E. 3.67

OpenStudy (anonymous):

The maximum possible score will be 4(3)+4(2)+4(4)+4(4)=52 The score earned was 2(3)+3(2)+4(4)+3(4)=40 A grade point of 4 equals the maximum score. So 40/52=x/4 So x= 3.0769 \[x \approx3.08\]